OpenStudy (anonymous):

Using the equation, 4Fe + 3O2 → 2Fe2O3, if 111.6 g of iron react, what mass of oxygen was consumed? 32.00 g 47.96 g 96.06 g 111.6 g

OpenStudy (abhisar):

Hi! From your equation you can see that 4 moles of iron (56*4=224 grams)need 3 moles of \(\sf O_2\) (32*3=96 grams), so for each gram of iron 96/224 = 0.428 grams of \(\sf O_2\) is needed. Now using unitary methods you can easily calculate amount of \(\sf O_2\) consumed for 111.6 grams of iron.

OpenStudy (abhisar):

Is it clear now?